# Date Fragments: A Hidden Bottleneck of Tokenization for Temporal Reasoning

### v1

@@ Abstract
Modern BPE tokenizers often split calendar dates into meaningless fragments, e.g., “20250312” $\rightarrow$ “202”, “503”, “12”, inflating token counts and obscuring the inherent structure needed for robust temporal reasoning. In this work, we (1) introduce a simple yet interpretable metric, termed date fragmentation ratio, that measures how faithfully a tokenizer preserves multi-digit date components; (2) release DateAugBench, a suite of 6500 examples spanning three temporal reasoning tasks: context-based date resolution, format-invariance puzzles, and date arithmetic across historical, contemporary, and future regimes; and (3) through layer-wise probing and causal attention-hop analyses, uncover an emergent date-abstraction mechanism whereby large language models stitch together the fragments of month, day, and year components for temporal reasoning. Our experiments show that excessive fragmentation correlates with accuracy drops of up to 10 points on uncommon dates like historical and futuristic dates. Further, we find that the larger the model, the faster the emergent date abstraction that heals date fragments is accomplished. Lastly, we observe a reasoning path that LLMs follow to assemble date fragments, typically differing from human interpretation (year $\rightarrow$ month $\rightarrow$ day).
Date Fragments: A Hidden Bottleneck of Tokenization
@@ 1 Introduction
Understanding and manipulating dates is a deceptively complex challenge for modern large language models (LLMs). Unlike ordinary words, dates combine numeric and lexical elements in rigidly defined patterns—ranging from compact eight-digit strings such as 20250314 to more verbose forms like “March 14, 2025” or locale-specific variants such as “14/03/2025.” Yet despite their structured nature, these date expressions often fall prey to subword tokenizers that fragment them into semantically meaningless pieces. A tokenizer that splits “2025-03-14” into “20”, “25”, “-0”, “3”, “-1”, “4” not only inflates the token count but also severs the natural boundaries of year, month, and day. This fragmentation obscures temporal cues and introduces a hidden bottleneck: even state-of-the-art LLMs struggle to resolve, compare, or compute dates accurately when their internal representations have been so badly fragmented. This issue is critical for real-world applications:
@@
Mis-tokenized dates can undermine scheduling and planning workflows, leading to erroneous calendar invites or appointments Vasileiou and Yeoh (2024). They can skew forecasting models in domains ranging from time-series analysis Tan et al. (2024); Chang et al. (2023) to temporal knowledge graph reasoning Wang et al. (2024). In digital humanities and historical scholarship, incorrect splitting of date expressions may corrupt timelines and misguide interpretative analyses Zeng (2024). As LLMs are increasingly deployed in cross-temporal applications, such as climate projectionWang and Karimi (2024), economic forecasting Carriero et al. (2024); Bhatia et al. (2024), and automated curriculum scheduling Vasileiou and Yeoh (2024), the brittleness introduced by subword fragmentation poses a risk of propagating temporal biases and inaccuracies into downstream scientific discoveries and decision-making systems Tan et al. (2024).
In this work, we provide a pioneer outlook on the impact of date tokenization on downstream temporal reasoning. Figure 1 illustrates how dates are processed internally for temporal reasoning.
Our contributions are summarized as follows:
@@
    We present date fragmentation ratio, a metric that measures how fragmented the tokenization outcome is compared to the actual year, month, and day components. We find that the fragmentation ratio generally correlates with temporal reasoning performance, namely that the more fragmented the tokenization, the worse the reasoning performance.
@@
    We analyse internal representations by tracing how LLMs “heal” fragmented date embeddings in their layer stack—an emergent ability that we term date abstraction. We find that larger models quickly can compensate for date fragmentation to achieve high accuracy of temporal reasoning at early layers.
@@
Our work fills the gap between tokenisation research Goldman et al. (2024); Schmidt et al. (2024) and temporal reasoning Su et al. (2024); Fatemi et al. (2024), and we suggest future work to consider date-aware vocabularies and adaptive tokenizers to ensure that date components remain intact.
@@ 2 Related Works
[…]
@@
Despite rapid progress on chain-of-thought and process-supervised reasoning, temporal cognition remains a conspicuous weakness of current LLMs. Benchmarks such as TimeBench Chu et al. (2024), TempReason Tan et al. (2023), Test-of-Time Fatemi et al. (2024), MenatQA Wei et al. (2023) and TimeQA Chen et al. (2021) reveal large gaps between model and human performance across ordering, arithmetic and co-temporal inference. Recent modelling efforts attack the problem from multiple angles: temporal-graph abstractions Xiong et al. (2024), instruction-tuned specialists such as Timo Su et al. (2024), pseudo-instruction augmentation for multi-hop QA Tan et al. (2023), and alignment techniques that re-ground pretrained models to specific calendar years Zhao et al. (2024). Yet these approaches assume a faithful internal representation of the input dates themselves. By introducing the notion of date fragmentation and demonstrating that heavier fragmentation predicts up to ten-point accuracy drops on DateAugBench, we uncover a failure mode that is orthogonal to reasoning algorithms or supervision: errors arise before the first transformer layer, at the level of subword segmentation. Addressing this front-end bottleneck complements, rather than competes with, existing efforts to improve temporal reasoning in LLMs.
@@ 3 DateAugBench
We introduce DateAugBench, benchmark designed to isolate the impact of date tokenisation on temporal reasoning in LLMs. DateAugBench comprises 6,500 augmented examples drawn from two established sources, TimeQA Chen et al. (2021) and TimeBench Chu et al. (2024), distributed across three tasks splits (see Table 1). Across all the splits, our chosen date formats cover a spectrum of common regional conventions (numeric with slashes, dashes, or dots; concatenated strings; two-digit versus four-digit years) and deliberately introduce fragmentation for atypical historical (e.g. “1799”) and future (e.g. “2121”) dates. This design enables controlled measurement of how tokenization compression ratios and subsequent embedding recovery influence temporal reasoning performance.
@@ 4 Experiment Design
4.1 Date Tokenization
Tokenizers.
For tokenization analysis, we compare a deterministic, rule-based baseline tokenizer against model-specific tokenizers. The baseline splits each date into its semantic components—year, month, day or Julian day—while preserving original delimiters. For neural models, we invoke either the OpenAI tiktoken encodings (for gpt-4, gpt-3.5-turbo, gpt-4o, text-davinci-003) or Hugging Face tokenizers for open-source checkpoints. Every date string is processed to record the resulting sub-tokens, token count, and reconstructed substrings.
@@ 4.1 Date Tokenization
Building on $\theta$, we introduce the date fragmentation ratio $F$, which quantifies how fragmented a tokenizer’s output is relative to the baseline. We initialise $F=0.0$ for a perfectly aligned segmentation and apply downward adjustments according to observed discrepancies: a 0.10 penalty if the actual year/month/day components are fragmented (i.e., $\mathbf{1}_{\mathrm{split}}=1$) , a 0.10 penalty if original delimiters are lost (i.e., $\mathbf{1}_{\mathrm{delimiter}}=1$), a 0.05 penalty multiplied by the token count difference $(N-N_{b}\bigr{)}$ between a tokenizer and the baseline, and a $0.30\times\theta$ penalty for distributional divergence. The resulting $F\in[0,1]$ provides an interpretable score: values close to 0 denote minimal fragmentation, and values near 1 indicate severe fragmentation.
@@
This date fragmentation ratio is pivotal because tokenisation inconsistencies directly impair a model’s ability to represent and reason over temporal inputs. When date strings are split non-intuitively, models face inflated token sequences and fragmented semantic cues, potentially leading to errors in tasks such as chronological comparison, date arithmetic, and context-based resolution. By quantifying fragmentation explicitly through $F$, we reveal hidden limitations in existing tokenizers, inform selections of robust architectures for time-sensitive applications.
@@ 4.2 Temporal Reasoning Evaluation
We evaluate a spectrum of model ranging from 0.5 B to 14 B parameters: five open-source Qwen 2.5 models (0.5 B, 1.5 B, 3 B, 7 B, 14 B) Yang et al. (2024), two Llama 3 models (3 B, 8 B) Touvron et al. (2023b), and two OLMo Groeneveld et al. (2024) models (1 B, 7 B). For comparison with state-of-the-art closed models, we also query the proprietary GPT-4o and GPT-4o-mini endpoints via the OpenAI API OpenAI et al. (2024).
@@
To measure how date tokenization affects downstream reasoning, we employ an LLM-as-judge framework using GPT-4o. For each test instance in DateFragBench, we construct a JSONL record that includes the question text, the model’s predicted answer, and a set of acceptable gold targets to capture all semantically equivalent date variants (e.g., both “03/04/2025” and “April 3, 2025” can appear in the gold label set). This record is submitted to GPT-4o via the OpenAI API with a system prompt instructing it to classify the prediction as CORRECT, INCORRECT, or NOT ATTEMPTED. A prediction is deemed CORRECT if it fully contains any one of the gold target variants without contradiction; INCORRECT if it contains factual errors relative to all gold variants; and NOT ATTEMPTED if it omits the required information. We validate GPT-4o’s reliability by randomly sampling 50 judged instances across all splits and obtaining independent annotations from four human reviewers. GPT-4o’s classifications agree with the human consensus on 97% of cases, yielding a Cohen’s $\kappa$ of 0.89, which affirms the reliability of our automated evaluation.
@@ 4.3 Internal Representations
We use four Qwen2.5 Yang et al. (2024) model checkpoints (0.5B, 1.5B, 3B, and 7B parameters) to trace how temporal information is processed internally across different layers. During inference, each question is prefixed with a fixed system prompt and a chain-of-thought cue, then passed through the model in evaluation mode. At each layer $i$, we extract the hidden-state vector corresponding to the final token position, yielding an embedding $h_{i}\in\mathbb{R}^{d}$ for that layer. Repeating over all examples produces a collection of layer-wise representations for positive and negative cases. We then quantify the emergence of temporal reasoning by training lightweight linear probes on these embeddings. For layer $i$, the probe is trained to distinguish “same-date” vs “different-date” examples. To explain when the model’s date understanding is achieved, we define the tokenization compensation point as the layer at which the model’s representation correctly represents the date in the given prompt. We experiment with this idea across various model sizes, aiming to test our hypothesis: larger models would recover calendar-level semantics from fragmented tokens at earlier stages, i.e., tokenization compensation is accomplished at early layers, as illustrated in Figure 2.
@@
We introduce a framework intended to understand in which order date fragments are stitched together for LLMs to answer a temporal question. Figure 1 depicts the idea of our framework: given an input prompt requiring a date resolution (e.g., “Is 28052025 the same date as 28th of May 2025?”), we define two sets of tokens: (1) concept tokens corresponding to year, month, and day fragments, and (2) decision tokens corresponding to the model answer (“yes” or “no”). Our framework aims to identify a stitching path for temporal reasoning, or reasoning path for short. A reasoning path is defined as a sequence of tokens containing date fragments and the model answer [^1] The idea of reasoning paths was introduced by Lindsey et al. (2025), which we leverage to interpret how LLMs address date fragments for temporal reasoning. . Given that there are multiple potential paths, we score each path and select the highest-scoring one as the LLM’s reasoning path for the given prompt. To score a reasoning path, our idea is the following: we identify when a date fragment or model answer is activated, by which input token and at which layer, and then determine how important each input token is for the date fragment and model answer. Our idea is implemented by using two different approaches: (i) next token prediction (§A.2.1): how likely a date fragment and model answer follows a given input token and (ii) token importance (§A.2.2): how important an input token is to a date fragment and model answer by replacing the input token with a random token. Lastly, we combine the results of the two approaches to yield the final score of a reasoning path (§A.2.3). This causal framework not only pinpoints where and when date fragments are activated, but also in which order they are stitched together to yield the model answer.
@@ 5.1 Date fragmentation
Table 2: Date fragmentation ratio across models and data splits over time. Here we refer to the family of models that use the same tokenizers.
@@
Table 3: Average accuracies per task. Context Rlt stands for context based resolution task, Fmt Switch refers to the format switching task, and Date Arth. refers to the date arithmetic task.
@@
Table 2 reports the mean date fragmentation ratio across four temporal regimes—Past (pre–2000), Near Past (2000–2009), Present (2010–2025), and Future (post–2025)—for each evaluated model. A ratio of 0.00 signifies perfect alignment with our rule-based baseline tokenizer, whereas higher values indicate progressively greater fragmentation. The rule-based Baseline unsurprisingly attains the maximal ratio of 0.00 in all periods, serving as a lower bound. Among neural architectures, OLMo Groeneveld et al. (2024) demonstrates the highest robustness, with an average fragmentation ratio of 0.15, closely followed by GPT-3 at 0.16. Both maintain strong fidelity across temporal splits, although performance dips modestly in the Future category (0.25), reflecting novel token sequences not seen during pre-training.
Model Tokenized output Frag-ratio
@@
A closer look, from Table 4, at sub-token granularity further explains these trends. Llama 3 Touvron et al. (2023b) and the GPT OpenAI et al. (2023) families typically segment each date component into three-digit sub-tokens (e.g., “202”, “504”, “03”), thus preserving the semantic unit of “MMDDYYYY” as compact pieces. OLMo Groeneveld et al. (2024) splits the date tokens into two digit tokens (e.g., “20”, “25”). By contrast, Qwen Yang et al. (2024) and Gemma Team et al. (2024) models break dates into single-digit tokens (e.g., “2”, “5”), whereas Phi Abdin et al. (2024) and LLama Touvron et al. (2023a) divide it into single-digit tokens with an initial token (e.g. “_”, “2”, “0”, “2”, “5”), inflating the token count. Although single-digit tokenisation can enhance models’ ability to perform arbitrary numeric manipulations (by treating each digit as an independent unit), it comes at the expense of temporal abstraction: the tight coupling between day, month, and year is lost, inflating the compression penalty and increasing the $\theta$ divergence from the baseline.

### v2

# Date Fragments: A Hidden Bottleneck of Tokenisation for Temporal Reasoning

@@ Abstract
Modern BPE tokenisers often split calendar dates into meaningless fragments, e.g., “20250312” $\rightarrow$ “202”, “503”, “12”, inflating token counts and obscuring the inherent structure needed for robust temporal reasoning. In this work, we (1) introduce a simple yet interpretable metric, termed date fragmentation ratio, that measures how faithfully a tokeniser preserves multi-digit date components; (2) release DateAugBench, a suite of 6500 examples spanning three temporal reasoning tasks: context-based date resolution, format-invariance puzzles, and date arithmetic across historical, contemporary, and future time periods; and (3) through layer-wise probing and causal attention-hop analyses, uncover an emergent date-abstraction mechanism whereby large language models stitch together the fragments of month, day, and year components for temporal reasoning. Our experiments show that excessive fragmentation correlates with accuracy drops of up to 10 points on uncommon dates like historical and futuristic dates. Further, we find that the larger the model, the faster the emergent date abstraction heals date fragments. Lastly, we observe a reasoning path that LLMs follow to assemble date fragments, typically differing from human interpretation (year $\rightarrow$ month $\rightarrow$ day). Our datasets and code are made publicly available here
Date Fragments: A Hidden Bottleneck of Tokenisation
@@ 1 Introduction
Understanding and manipulating dates is a deceptively complex challenge for modern large language models (LLMs). Unlike ordinary words, dates combine numeric and lexical elements in rigidly defined patterns—ranging from compact eight-digit strings such as 20250314 to more verbose forms like “March 14, 2025” or locale-specific variants such as “14/03/2025.” Yet despite their structured nature, these date expressions often fall prey to subword tokenisers that fragment them into semantically meaningless pieces. A tokeniser that splits “2025-03-14” into “20”, “25”, “-0”, “3”, “-1”, “4” not only inflates the token count but also severs the natural boundaries of year, month, and day. This fragmentation obscures temporal cues and introduces a hidden bottleneck: even state-of-the-art LLMs struggle to resolve, compare, or compute dates accurately when their internal representations have been so badly fragmented. This issue has a critical impact on real-world applications:
@@
Mis-tokenised dates can undermine scheduling and planning workflows, leading to erroneous calendar invites or appointments Vasileiou and Yeoh (2024). They can skew forecasting models in domains ranging from time-series analysis Tan et al. (2024); Chang et al. (2023) to temporal knowledge graph reasoning Wang et al. (2024). In digital humanities and historical scholarship, incorrect splitting of date expressions may corrupt timelines and misguide interpretative analyses Zeng (2024). As LLMs are increasingly deployed in cross-temporal applications, such as climate projection Wang and Karimi (2024), economic forecasting Carriero et al. (2024); Bhatia et al. (2024), and automated curriculum scheduling Vasileiou and Yeoh (2024), the brittleness introduced by subword fragmentation poses a risk of propagating temporal biases and inaccuracies into downstream scientific discoveries and decision-making systems Tan et al. (2024).
In this work, we provide a pioneer outlook on the impact of date tokenisation on downstream temporal reasoning. Figure 1 illustrates how dates are processed internally for temporal reasoning. Our contributions are summarized as follows:
@@
    We present date fragmentation ratio, a metric that measures how fragmented the tokenisation outcome is compared to the actual year, month, and day components. We find that the fragmentation ratio generally correlates with temporal reasoning performance, namely that the more fragmented the tokenisation, the worse the reasoning performance.
@@
    We analyse internal representations by tracing how LLMs “heal” a fragmented date embeddings in their layer stack—an emergent ability that we term date abstraction. We find that larger models can quickly compensate for date fragmentation at early layers to achieve high accuracy for date equivalence reasoning.
@@
Our work fills the gap between tokenisation research Goldman et al. (2024); Schmidt et al. (2024) and temporal reasoning Su et al. (2024); Fatemi et al. (2024), and we suggest future work to consider date-aware vocabularies and adaptive tokenisers to ensure that date components remain intact.
@@ 2 Related Works
Recent scholarship interrogates four complementary facets of sub-word segmentation: (i) tokenisation fidelity, i.e. how closely a tokeniser preserves semantic units: Large empirical studies show that higher compression fidelity predicts better downstream accuracy in symbol-heavy domains such as code, maths and dates Goldman et al. (2024); Schmidt et al. (2024); (ii) numeric segmentation strategies that decide between digit-level or multi-digit units: Previous work demonstrates that the choice of radix-single digits versus 1-3 digit chunks induces stereotyped arithmetic errors and can even alter the complexity class of the computations LLMs can realise Singh and Strouse (2024); Zhou et al. (2024); (iii) probabilistic or learnable tokenisers whose segmentations are optimised jointly with the language model: Theory frames tokenisation as a stochastic map whose invertibility controls whether maximum-likelihood estimators over tokens are consistent with the underlying word distribution Gastaldi et al. (2024); Rajaraman et al. (2024) and (iv) pre-/post-tokenisation adaptations that retrofit a model with a new vocabulary: Zheng et al. (2024) introduce an adaptive tokeniser that co-evolves with the language model, while Liu et al. (2025) push beyond the “sub-word” dogma with SuperBPE, a curriculum that first learns subwords and then merges them into cross-whitespace “superwords”, cutting average sequence length by 27 %. Complementary studies expose and correct systematic biases introduced by segmentation Phan et al. (2024) and propose trans-tokenisation to transfer vocabularies across languages without re-training the model from scratch Remy et al. (2024). Our work builds on these insights but zooms in on calendar dates—a hybrid of digits and lexical delimiters whose multi-digit fields are routinely shredded by standard BPE, obscuring cross-field regularities crucial for temporal reasoning.
@@
Despite rapid progress on chain-of-thought and process-supervised reasoning, temporal cognition remains a conspicuous weakness of current LLMs. Benchmarks such as TimeBench Chu et al. (2024), TempReason Tan et al. (2023), Test-of-Time Fatemi et al. (2024), MenatQA Wei et al. (2023) and TimeQA Chen et al. (2021) reveal large gaps between model and human performance across ordering, arithmetic and co-temporal inference. Recent modelling efforts attack the problem from multiple angles: temporal-graph abstractions Xiong et al. (2024), instruction-tuned specialists such as Timo Su et al. (2024), pseudo-instruction augmentation for multi-hop QA Tan et al. (2023), and alignment techniques that re-ground pretrained models to specific calendar years Zhao et al. (2024). Yet these approaches assume a faithful internal representation of the input dates themselves. By introducing the notion of date fragmentation and demonstrating that heavier fragmentation predicts up to ten-point accuracy drops on DateAugBench, we uncover a failure mode that is orthogonal to reasoning algorithms or supervision: errors arise before the first transformer layer, at the level of subword segmentation. Addressing this front-end bottleneck complements existing efforts to further improve LLMs for temporal reasoning.
@@ 3 DateAugBench
We introduce DateAugBench, benchmark designed to isolate the impact of date tokenisation on temporal reasoning in LLMs. DateAugBench comprises 6,500 augmented examples drawn from two established sources, TimeQA Chen et al. (2021) and TimeBench Chu et al. (2024), distributed across three tasks splits (see Table 1). Across all the splits, our chosen date formats cover a spectrum of common regional conventions (numeric with slashes, dashes, or dots; concatenated strings; two-digit versus four-digit years) and deliberately introduce fragmentation for atypical historical (e.g. “1799”) and future (e.g. “2121”) dates. This design enables controlled measurement of how tokenisation compression ratios and subsequent embedding recovery influence temporal reasoning performance.
@@ 4 Experiment Design
4.1 Date Tokenisation
Tokenisers.
For tokenisation analysis, we compare a deterministic, rule-based baseline tokeniser against model-specific tokenisers. The baseline splits each date into its semantic components—year, month, day or Julian day—while preserving original delimiters. For neural models, we invoke either the OpenAI TikTok encodings (for gpt-4, gpt-3.5-turbo, gpt-4o, text-davinci-003) or Hugging Face tokenisers for open-source checkpoints. Every date string is processed to record the resulting sub-tokens, token count, and reconstructed substrings.
@@ 4.1 Date Tokenisation
Building on $\theta$, we introduce the date fragmentation ratio $F$, which quantifies how fragmented a tokeniser’s output is relative to the baseline. We initialise $F=0.0$ for a perfectly aligned segmentation and apply downward adjustments according to observed discrepancies: a 0.10 penalty if the actual year/month/day components are fragmented (i.e., $\mathbf{1}_{\mathrm{split}}=1$) , a 0.10 penalty if original delimiters are lost (i.e., $\mathbf{1}_{\mathrm{delimiter}}=1$), a 0.05 penalty multiplied by the token count difference $(N-N_{b}\bigr{)}$ between a tokeniser and the baseline, and a $0.30\times\theta$ penalty for distributional divergence. The resulting $F\in[0,1]$ provides an interpretable score: values close to 0 denote minimal fragmentation, and values near 1 indicate severe fragmentation.
@@
This date fragmentation ratio is pivotal because tokenisation inconsistencies directly impair a model’s ability to represent and reason over temporal inputs. When date strings are split non-intuitively, models face inflated token sequences and fragmented semantic cues, potentially leading to errors in tasks such as chronological comparison, date arithmetic, and context-based resolution.
@@ 4.2 Temporal Reasoning Evaluation
We evaluate a spectrum of model ranging from 0.5 B to 14 B parameters: five open-source Qwen 2.5 models (0.5 B, 1.5 B, 3 B, 7 B, 14 B) Yang et al. (2024), two Llama 3 models (3 B, 8 B) Touvron et al. (2023), and two OLMo Groeneveld et al. (2024) models (1 B, 7 B). For comparison with state-of-the-art closed models, we also query the proprietary GPT-4o and GPT-4o-mini endpoints via the OpenAI API OpenAI et al. (2024).
@@
To measure how date tokenisation affects downstream reasoning, we employ an LLM-as-judge framework using GPT-4o. For each test instance in DateFragBench, we construct a JSONL record that includes the question text, the model’s predicted answer, and a set of acceptable gold targets to capture all semantically equivalent date variants (e.g., both “03/04/2025” and “April 3, 2025” can appear in the gold label set). This record is submitted to GPT-4o via the OpenAI API with a system prompt instructing it to classify the prediction as CORRECT, INCORRECT, or NOT ATTEMPTED. A prediction is deemed CORRECT if it fully contains any one of the gold target variants without contradiction; INCORRECT if it contains factual errors relative to all gold variants; and NOT ATTEMPTED if it omits the required information. We validate GPT-4o’s reliability by randomly sampling 50 judged instances across all splits and obtaining independent annotations from four student evaluators. In 97% of cases, GPT-4o’s judgments of model answers agree with the averaged human judgments across four student evaluators, with a Cohen’s $\kappa$ of 0.89 as the inter-annotator agreement, affirming the reliability of our automatic evaluation setup.
@@ 4.3 Internal Representations
We use four Qwen2.5 Yang et al. (2024) model checkpoints (0.5B, 1.5B, 3B, and 7B parameters) to trace how temporal information is processed internally across different layers. During inference, each question is prefixed with a fixed system prompt and a chain-of-thought cue, then passed through the model in evaluation mode. At each layer $i$, we extract the hidden-state vector corresponding to the final token position, yielding an embedding $h_{i}\in\mathbb{R}^{d}$ for that layer. Repeating over all examples produces a collection of layer-wise representations for positive and negative cases. We then quantify the emergence of temporal reasoning by training lightweight linear probes on these embeddings. For layer $i$, the probe is trained to distinguish “same-date” (positive) vs “different-date” (negative) examples. To explain when the model’s date understanding is achieved, we define the tokenisation compensation point as the layer at which the model’s representation correctly represents the date in the given prompt. We experiment with this idea across various model sizes, aiming to test our hypothesis: larger models would recover calendar-level semantics from fragmented tokens at earlier stages, i.e., tokenisation compensation is accomplished at early layers, as illustrated in Figure 2.
@@
We introduce a framework intended to understand in which order date fragments are stitched together for LLMs to answer a temporal question. Figure 1 depicts the idea of our framework: given an input prompt requiring a date resolution (e.g., “Is 28052025 the same date as 28th of May 2025?”), we define two sets of tokens: (1) concept tokens corresponding to year, month, and day fragments, and (2) decision tokens corresponding to the model answer (“yes” or “no”). Our framework aims to identify a stitching path for temporal reasoning, or reasoning path for short. A reasoning path is defined as a sequence of tokens containing date fragments and the model answer [^1] The idea of reasoning paths was introduced by Lindsey et al. (2025), which we leverage to interpret how LLMs address date fragments for temporal reasoning. . Given that there are multiple potential paths, we score each path and select the highest-scoring one as the LLM’s reasoning path for the given prompt. To score a reasoning path, our idea is the following: we identify when a date fragment or model answer is activated, by which input token and at which layer, and then determine how important each input token is for the date fragment and model answer. Our idea is implemented by using two different approaches: (i) next token prediction (§A.2.1): how likely a date fragment and model answer follows a given input token and (ii) token importance (§A.2.2): how important an input token is to a date fragment and model answer (by replacing the input token with a random token). Lastly, we combine the results of the two approaches to yield the final score of a reasoning path (§A.2.3). This causal framework not only pinpoints where and when date fragments are activated, but also in which order they are stitched together to yield the model answer.
@@ 5.1 Date fragmentation
Table 2: Date fragmentation ratio across models and data splits over time. In case a family of model variants (Qwen, Gemma, DeepSeek and Phi) uses the same tokeniser, only the family name is referenced.
@@
Table 3: Average accuracies per task. Context Rlt stands for context based resolution, Fmt Switch refers to format switching, and Date Arth. refers to date arithmetic.
@@
Table 2 reports the mean date fragmentation ratio across four time periodsPast (pre–2000), Near Past (2000–2009), Present (2010–2025), and Future (post–2025)—for each evaluated model. A ratio of 0.00 signifies perfect alignment with our rule-based baseline tokeniser, whereas higher values indicate progressively greater fragmentation. The rule-based Baseline unsurprisingly attains the maximal ratio of 0.00 in all periods, serving as a lower bound. Among neural architectures, OLMo Groeneveld et al. (2024) demonstrates the highest robustness, with an average fragmentation ratio of 0.15, closely followed by GPT-3 at 0.16. Both maintain strong fidelity across temporal splits, although performance dips modestly in the Future category (0.25), reflecting novel token sequences not seen during pre-training.
Model Tokenised output Frag-ratio
@@
A closer look, from Table 4, at sub-token granularity further explains these trends. Llama 3 Touvron et al. (2023) and the GPT OpenAI et al. (2023) families typically segment each date component into three-digit sub-tokens (e.g., “202”, “504”, “03”), thus preserving the semantic unit of “MMDDYYYY” as compact pieces. OLMo Groeneveld et al. (2024) splits the date tokens into two digit tokens (e.g., “20”, “25”). By contrast, Qwen Yang et al. (2024) and Gemma Team et al. (2024) models break dates into single-digit tokens (e.g., “2”, “5”), whereas Phi Abdin et al. (2024) divides it into single-digit tokens with an initial token (e.g. “_”, “2”, “0”, “2”, “5”), inflating the token count. Although single-digit tokenisation can enhance models’ ability to perform arbitrary numeric manipulations (by treating each digit as an independent unit), it comes at the expense of temporal abstraction: the tight coupling between day, month, and year is lost, inflating the compression penalty and increasing the $\theta$ divergence from the baseline.